Details.

10:10am

Power on Ok.

OCAM noise quadrants.

TN:20230204_173447
  • Power cycle BCU2k
  • rebin
  • OK
Preset AO.

Sat Feb 4 17:50:28.501 2023 aos.PresetAO{FLAO}.started - PresetAO {FLAO}[780] (aomode:ACE-AO,focst:bentGregorianFront,instr:LUCIFER,wfs:FLAOWFS,sox:0[0],soy:0,rox:0[0],roy:0,mag:10,cindex:0,,elev:nan,rotang:nan,grvang:nan,r0:0,skybr:nan,windsp:nan,winddir:nan)

Observed a group of actuators not force control, positions seems correct.

TN:_20230204_182904

cancel preset.

18:29:56.016752| MAIN > Maximum mode index used to apply flat -> # 605

close loop.

Gain0.1, observed the spots at 4oclock. Act 498,501,533,502

adsec_act20230204.png

saved TN:_184128

Talk with Guido about it. he will be available during the afternoon.

I decided not continue the AO pupil monitor and check.

11:30am

Adsec rest

WFS set safe.

3:00PM contacted Guido and talk about t. he will analyse it and talk with Xianyu if it require to remove actuator.

-- JuanCarlosGuerra - 04 Feb 2023
Topic revision: r3 - 05 Feb 2023, JuanCarlosGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback